OpenStudy (anonymous):

Want to double check on this: 3(x + 5) - 2(x + 2) = 20 I got 1 as the answer. Is that right?

OpenStudy (anonymous):

na. The answer should be x=9

OpenStudy (anonymous):

If you doubt your answer then put 1 in place of x in your equation and check whether you are getting both sides equal or not...

OpenStudy (anonymous):

3(1 + 5) - 2(1 + 2) = 20 18 - 6 = 20 12 = 20 ( false) This means 1 is not the right answer..

OpenStudy (anonymous):

Darn! do you know what I might have done wrong?

OpenStudy (anonymous):

Let us go step by step...

OpenStudy (anonymous):

Distribute 3 and 2 to their respective brackets first... Can you do that??

OpenStudy (anonymous):

\(3(x + 5) = ??\)

OpenStudy (anonymous):

You could have made the mistake when you expanded the brackets... For example -2(x+2) is -2x-4

OpenStudy (anonymous):

Wow! yeah I did...I didn't think it was going to be negative.

OpenStudy (anonymous):

3x+15

OpenStudy (anonymous):

\(-2(x + 2) = ??\)

OpenStudy (anonymous):

-2x+4?

OpenStudy (anonymous):

Nope here you have mistaken..

OpenStudy (anonymous):

always look out for a negative if the number which you are multyplying the bracket with is negative. Do not leave out the negative sign

OpenStudy (anonymous):

\[-2 ( x + 2) \implies -2 \times x + (-2) \times (2) \implies -2x + (-4) \implies -2x - 4\]

OpenStudy (anonymous):

So that is where you went wrong @jewjewbri in -2(x+2)

OpenStudy (anonymous):

Remember : \(+(-) = \; \; -\)